OpenStudy (raffle_snaffle):

Let H be the set of points inside and the unit circle in the xy-plane. THat is, let H={[x y] |x^2+y^2<=1}. Find a specific example--two vectors or a vector and a scalar-to show that H is not a subspace of R^2.

OpenStudy (raffle_snaffle):

@Loser66

OpenStudy (raffle_snaffle):

I let v=[1/root(2) ,1/root(2)]

OpenStudy (raffle_snaffle):

then v+v=[1/root(2) ,1/root(2)]+[1/root(2) ,1/root(2)]

OpenStudy (raffle_snaffle):

solution will be v=[root(2) root(2)]

OpenStudy (raffle_snaffle):

that is suppose to be sqrt

OpenStudy (loser66):

let u = (x, y) \(\in H\) so, x^2+y^2 =1 let v=(z, t) \(\in H \) so that z^2 + t^2 =1 so, u + v is supposed to be in H, right? now see, u+v = x^2 +y^2 +z^2+t^2 =4 ( not \(\leq 1) \), therefore, H is not a subspace of R^2

OpenStudy (loser66):

sorry, =2, hehehe

OpenStudy (raffle_snaffle):

how is u+v in H?

OpenStudy (loser66):

by definition of the subspace , V is a vector space, (this case is R^2) H is a set in V , H is a subspace when : 1/ u in H, v in H , then u + v in H 2/ there is a scalar a such that u in H , a*u in H If it satisfies 2 conditions above, then H is a subspace. Your problem is prove it is NOT a subspace, so, just pick the easiest way to show that it is NOT satisfy either of the conditions. I pick =1 to easy compute. if you pick < 1, it's hard to prove that the sum is >1 , fortunately, your problem give us the case =1. so?? take advantage on it. hehehe

OpenStudy (loser66):

woaah, Zarkon is here. let him give us his opinion. I dare not to say anything in the front of him. In other word, I am intimidated by hiiiiiim, hhehehehe.

OpenStudy (raffle_snaffle):

That helped out a lot.

OpenStudy (zarkon):

you could also use... \[(1/2,0)\in H\] but \[4(1/2,0)=(2,0)\notin H\]

OpenStudy (raffle_snaffle):

How does u+v = x^2 +y^2 +z^2+t^2 =4 ( not ≤1)

OpenStudy (loser66):

I correct it, it is =2 , not 4 ( I am sorry for that) hey, !!! how 2 \(\leq \)1?

OpenStudy (raffle_snaffle):

why did you say x^2 +y^2 +z^2+t^2 why did you include z and t that would make R^4

OpenStudy (loser66):

2 is NOT \(\leq \)1 , right? so, u + v is not in H, right?

OpenStudy (raffle_snaffle):

Yes I agree. Lol

OpenStudy (loser66):

x,y is 2 elements of vector u in H so that x^2 + y^2 =1 , Ok, take another circle to make it different v = x^2 + y^2 = 0.9

OpenStudy (loser66):



OpenStudy (raffle_snaffle):

^Yes yes I agree.

OpenStudy (loser66):